# Schur–Weyl Equivalences for Wreath Product Superalgebras

### v1

@@ Abstract.
Let $A$ be an associative superalgebra over a field of characteristic zero. Let $n\geq d+1$. The main result of the paper establishes an equivalence of categories between supermodules for the wreath product $S_{d}\wr A$ and an explicitly defined category of supermodules for the general linear Lie algebra $\mathfrak{gl}_{n}(A)$. We also give an example showing the bound $n\geq d+1$ cannot be improved.
@@ 1.1. Overview
Schur–Weyl duality has played a fundamental role in representation theory for more than century. It provides an equivalence of categories between the complex representations of the symmetric group on $d$ letters, $S_{d}$, and the polynomial representations of degree $d$ for the general linear Lie algebra, $\mathfrak{gl}_{n}(\mathbb{C})$, whenever $n\geq d$. Due to its importance, Schur–Weyl duality has been generalized to a wide variety of settings. Relevant to this paper, these generalizations include when $\mathfrak{gl}_{n}(\mathbb{C})$ is replaced with affine and toroidal Lie algebras, or with the type $Q$ Lie superalgebra [13, 7, 28]. In which case, the symmetric group is replaced with the affine symmetric group or Sergeev superalgebra, respectively. The purpose of the present paper is to provide a uniform proof that includes all of these, along with many additional cases of interest.
@@ 1.2. Main Result
For $d\geq 1$, let $S_{d}\wr A$ denote the wreath product superalgebra of $\mathds{k}S_{d}$ and $A$. Wreath product superalgebras naturally generalize many of the nice features of the symmetric groups and have long been of interest. On the Lie algebra side, for $n\geq 1$, let $\mathfrak{gl}_{n}(A)$ denote the Lie superalgebra over $\mathds{k}$ consisting of $n\times n$-matrices with entries from $A$ and Lie bracket given by the supercommutator. This also includes many examples of interest; see Section 2.3. We refer the reader to Sections 2.2 and 2.5 for details on these two superalgebras.
@@
In Definition 2.4.1 we define a full subcategory, $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$, of $U(\mathfrak{gl}_{n}(A))$-supermodules consisting of supermodules which are polynomial representations of degree $d$ when restricted to the subalgebra $\mathfrak{gl}_{n}(\mathds{k})\subseteq\mathfrak{gl}_{n}(A)$ and also satisfy a mild technical condition. The image of $F^{A}_{n,d}$ is easily seen to land in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. Our main theorem is the following result.
@@
When $A=\mathds{k}[t_{1}^{\pm 1},\dotsc,t_{r}^{\pm 1}]$ or $A$ is the Clifford superalgebra on one generator, then we essentially recover the aforementioned generalizations of classical Schur–Weyl duality. However, here $A$ is an arbitrary associative $\mathds{k}$-superalgebra. We also note that Example 3.8.1 shows that the bound $n\geq d+1$ is sharp. Thus, in some sense, this is the most general result along these lines that one could hope for.
Our general approach is similar to that of [13, 7]: we prove the equivalence by bootstrapping from the classical Schur–Weyl duality equivalence. This paper is an instance of George Pólya’s Inventor’s Paradox: it turns out that our more general situation makes for easier proofs. As far as we are aware, the technique of building from the classical case goes back to Chari–Pressley [9].
@@ 1.3. Future Directions
There is a Schur–Weyl equivalence between the Lie superalgebra $\mathfrak{gl}_{m|n}(\mathbb{C})$ and the symmetric group [5, 28]. We fully expect the methods used here should generalize to this case and yield an equivalence of categories between certain $\mathfrak{gl}_{m|n}(A)$-supermodules and $S_{d}\wr A$-supermodules. See [14] where something like this is done for $A=\mathds{k}[t^{\pm 1}]$.
Classical Schur–Weyl duality holds for any infinite field, and one could hope to prove a positive characteristic analogue of the main theorem. Properly formulating a definition of the category $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$ will likely require working with an appropriate integral form for $U(\mathfrak{gl}_{n}(A))$. The existence of integral forms for the universal enveloping superalgebras of affine and map Lie (super)algebras suggests this should be possible [1, 8, 15]. The representation theory of quantized and/or affinized wreath product superalgebras and their associated quantized Schur algebras is likewise interesting. See [21, 26, 27, 30] and the references therein. It would be interesting to obtain a quantized and/or affinized version of the main theorem. See [9] for the quantized affine case.
@@ 2.1. Notations and Conventions
$I(n,d)=\left\{\mathbf{i}=(i_{1},\dotsc,i_{d})\mid i_{1},\dotsc,i_{d}\in[1,n]\right\}.$
Let $\tilde{I}(n,d)\subseteq I(n,d)$ be the subset of tuples that have pairwise distinct entries. Let
@@ 2.2. Lie Superalgebras
The set of matrices $\{E_{r,s}^{b}\mid r,s\in[1,n]\text{ and }b\in\mathcal{B}\}$ form a homogeneous $\mathds{k}$-basis for $\mathfrak{gl}_{n}(A)$. We view $\mathfrak{gl}_{n}(A)$ as a Lie superalgebra over $\mathds{k}$ via the graded commutator bracket,
$[E_{r,s}^{a},E_{r^{\prime},s^{\prime}}^{a^{\prime}}]=\delta_{s,r^{\prime}}E_{r,s^{\prime}}^{aa^{\prime}}-(-1)^{\overline{a}\cdot\overline{a^{\prime}}}\delta_{s^{\prime},r}E_{r^{\prime},s}^{a^{\prime}a},$
for all $r,s,r^{\prime},s^{\prime}\in[1,n]$ and homogeneous $a,a^{\prime}\in A$. As we did here, we often only give definitions and formulas on homogeneous elements. The general case is obtained by extending linearly.
@@
For short, write $h_{i}=E^{1}_{i,i}$ for $i=1,\dotsc,n$, and let $\mathfrak{h}$ denote the Cartan subalgebra of $\mathfrak{gl}_{n}(\mathds{k})$ with basis $h_{1},\dotsc,h_{n}$. Let $\varepsilon_{i}\in\mathfrak{h}^{*}=\operatorname{Hom}_{\mathds{k}}(\mathfrak{h},\mathds{k})$ be the linear functional determined by $\varepsilon_{i}(h_{j})=\delta_{i,j}$. Let $\Lambda(n)=\oplus_{i=1}^{n}\mathbb{Z}\varepsilon_{i}$ and let $\Lambda^{+}(n)=\oplus_{i=i}^{n}\mathbb{Z}_{\geq 0}\varepsilon_{i}$. For $d\geq 0$, let
$\Lambda^{+}(n,d)=\left\{\lambda=\sum_{i=1}^{n}\lambda_{i}\varepsilon_{i}\in\Lambda^{+}(n)\mid\sum_{i=1}^{n}\lambda_{i}=d\right\}.$
For short, given $\mathbf{i}=(i_{1},\dotsc,i_{d})\in I(n,d)$, write $\boldsymbol{\varepsilon}_{\mathbf{i}}=\sum_{k=1}^{d}\varepsilon_{i_{k}}\in\Lambda^{+}(n,d)$.
@@ 2.3. Examples
The above encompasses a number of cases of interest. We give a few examples.
If $A=\mathds{k}$, then $\mathfrak{gl}_{n}(\mathds{k})$ is the general linear Lie algebra. If $A=\operatorname{Cliff}:=\mathds{k}[c]/(c^{2}-1)$ is the Clifford superalgebra on one odd generator, then $\mathfrak{gl}_{n}(A)\cong\mathfrak{q}_{n}(\mathds{k})$ is the type Q Lie superalgebra. When $\Lambda=\mathds{k}[x]/(x^{2})$ is the Grassman algebra on one odd generator, the Lie superalgebra $\mathfrak{sl}_{2}(\Lambda)$ and its representations appear in the context of annular Khovanov homology [17].
If $A=\mathds{k}[t_{1},\dotsc,t_{k}]$ or $A=\mathds{k}[t_{1}^{\pm 1},\dotsc,t_{k}^{\pm 1}]$, then $\mathfrak{gl}_{n}(A)$ is the general linear (multi)current Lie algebra or (multi)loop Lie algebra, respectively. The representations of (multi)loop Lie algebras are closely related to the level zero representations of the general linear affine and toroidal Lie algebras. A noncommutative example of this kind is when $A=\mathds{k}_{q}$, a quantum torus. In this case, $\mathfrak{sl}_{n}(\mathds{k}_{q})$ appears as the core of the elliptic quasi-simple Lie algebras in type A [2]. Current algebras and their representations are central to the study of Yangians and, more recently, have appeared in the trace decategorifications of categorified quantum groups [23, 3, 29]. Generalizing the loop and current Lie algebras, if $A$ is a finitely-generated commutative $\mathds{k}$-algebra, then $\mathfrak{gl}_{n}(A)$ is a so-called map Lie algebra and its representations are an area of active research.
Given a superalgebra $A$ and a superalgebra involution $\iota:A\to A$, one can define an associated superalgebra $\operatorname{Cliff}(A)$, which equals $A\otimes\operatorname{Cliff}$ as a superspace and has product given by $(a_{1}\otimes c^{k_{1}})(a_{2}\otimes c^{k_{2}})=(-1)^{\overline{a_{2}}\cdot\overline{k_{1}}}a_{1}\iota^{k_{1}}(a_{2})\otimes c^{k_{1}+k_{2}}$ for homogeneous $a_{1},a_{2}\in A$ and $k_{1},k_{2}\in\{0,1\}$. Then one could define
$\mathfrak{q}_{n}(A)=\mathfrak{gl}_{n}(A\otimes\operatorname{Cliff}).$
Taking $A=\mathds{k}$ and $\iota$ to be the identity map recovers the case $\mathfrak{q}_{n}(\mathds{k})\cong\mathfrak{gl}_{n}(\operatorname{Cliff})$ mentioned in the previous paragraph. Taking $A=\mathds{k}[t]$ or $A=\mathds{k}[t^{\pm 1}]$ and $\iota:A\to A$ determined by $t\mapsto-t$, or $A=\mathds{k}[t^{\pm 1}]$ and $\iota:A\to A$ determined by $t\mapsto t^{-1}$ recovers the twisted Lie superalgebras of type Q studied by Nazarov [24] and Chen–Guay [6]. The Lie superalgebra $\mathfrak{gl}_{n}(A)$ with $A=\mathds{k}\langle u^{\pm 1},c\rangle/(uu^{-1}=u^{-1}u=1,c^{2}=u,cu=uc)$ with $u$ even and $c$ odd is closely related to the loop algebra for the twisted affine type Q Lie superalgebra $\mathfrak{q}(n)^{(2)}$ studied in, for example, [20, 19].
@@ 2.4. Representations
A $\mathfrak{gl}_{n}(A)$-supermodule is a $\mathds{k}$-superspace $M=M_{\bar{0}}\oplus M_{\bar{1}}$ which is a $U(\mathfrak{gl}_{n}(A))$-module where the action respects the parity in that $x.m\in M_{\overline{x}+\overline{m}}$ for all homogeneous $x\in U(\mathfrak{gl}_{n}(A))$ and $m\in M$. There is a natural $\mathbb{Z}_{2}$-grading on $\operatorname{Hom}_{\mathds{k}}(M,N)$ where $f:M\to N$ is declared to have parity $r\in\mathbb{Z}_{2}$ if $f(M_{s})\subseteq N_{s+r}$ for all $s\in\mathbb{Z}_{2}$. A homogeneous $\mathfrak{gl}_{n}(A)$-module homomorphism is a homogeneous map $f:M\to N$ which satisfies $f(x.m)=(-1)^{\overline{f}\cdot\overline{x}}f(m)$ for all homogeneous $x\in U(\mathfrak{gl}_{n}(A))$ and $m\in M$. A $\mathfrak{gl}_{n}(A)$-supermodule homomorphism is a $\mathds{k}$-linear map $f:M\to N$ that can be written $f=f_{\bar{0}}+f_{\bar{1}}$ where $f_{r}$ is a homogeneous $\mathfrak{gl}_{n}(A)$-supermodule homomorphism of parity $r\in\mathbb{Z}_{2}$. In particular, note that we do not assume supermodule homomorphisms preserve parity.
@@
for all $i=1,\dotsc,n$. We write $M_{\lambda}$ for the subspace of all weight vectors of weight $\lambda$, and call $M$ a weight module if $M=\oplus_{\lambda\in\mathfrak{h}^{*}}M_{\lambda}$. The supercommutator formula implies that
@@
Set $V_{n}=A^{\oplus n}$ to be column vectors of height $n$ with entries from $A$. This is naturally a $\mathfrak{gl}_{n}(A)$-supermodule with $\mathbb{Z}_{2}$-grading given by $V_{n,r}=A_{r}^{\oplus n}$ for $r\in\mathbb{Z}_{2}$, and with action given by matrix multiplication. Since $U(\mathfrak{gl}_{n}(A))$ is a Hopf superalgebra there is a natural action on $V_{n}^{\otimes d}$ for all $d\geq 1$. For $i\in[1,n]$ and homogeneous $a\in A$, write $v_{i}^{a}$ for a column vector with $a$ in the $i$th row, and zeros elsewhere. The parity of $v_{i}^{a}$ is $\overline{a}$. Using this notation,
@@
is a homogeneous $\mathds{k}$-basis for $V_{n}^{\otimes d}$. From this it is easy to see that $V_{n}^{\otimes d}$ is a weight module, $v_{i_{1}}^{b_{1}}\otimes\dotsb\otimes v_{i_{d}}^{b_{d}}\in\left(V_{n}^{\otimes d}\right)_{\boldsymbol{\varepsilon}_{\mathbf{i}}}$, and that the weights of $V_{n}^{\otimes d}$ lie in $\Lambda^{+}(n,d)$.
Write $E_{n}=\mathds{k}^{\oplus n}$ for the natural representation of $\mathfrak{gl}_{n}(\mathds{k})$; that is, $E_{n}$ consists of column vectors of height $n$ where $\mathfrak{gl}_{n}(\mathds{k})$ acts by matrix multiplication. Recall that for $d\geq 0$, the irreducible polynomial representations of degree $d$ of $\mathfrak{gl}_{n}(\mathds{k})$ are, by definition, the irreducible representations which appear as a composition factor of $E_{n}^{\otimes d}$. We follow the convention that $V_{n}^{\otimes 0}=E_{n}^{\otimes 0}=\mathds{k}$ is the trivial module.
For the purposes of this paper, a polynomial representation of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ will mean a $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodule which is isomorphic to a (possibly infinite) direct sum of irreducible polynomial representations of degree $d$. The full subcategory of polynomial representations of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ is a semisimple category. If $M$ is a polynomial representation of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$, then $M$ is a weight module with
$M=\bigoplus_{\lambda\in\Lambda^{+}(n,d)}M_{\lambda}.$
@@
To foreshadow what comes later, $U(\mathfrak{gl}_{n}(A))$-supermodules of the form $V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N$, where the action of $U(\mathfrak{gl}_{n}(A))$ is on $V_{n}^{\otimes d}$ also satisfy both of the above conditions. Nevertheless, perhaps the second condition could use some elaboration. We note it already appears in the study of polynomial representations of $\mathfrak{q}_{n}(\mathds{k})=\mathfrak{gl}_{n}(\operatorname{Cliff})$ [16, Definition 1.5]. The following example shows that this condition is not superfluous.

### v2

@@ Abstract.
Let $A$ be an associative superalgebra over a field of characteristic zero. Let $n\geq d+1$. The main result of the paper establishes an equivalence between the category of supermodules for the wreath product $S_{d}\wr A$ and an explicitly defined category of supermodules for the general linear Lie superalgebra $\mathfrak{gl}_{n}(A)$. We also give an example showing the bound $n\geq d+1$ cannot be improved.
@@ 1.1. Overview
Schur–Weyl duality has played a fundamental role in representation theory for more than century. It provides an equivalence of categories between the complex representations of the symmetric group on $d$ letters, $S_{d}$, and the polynomial representations of degree $d$ for the general linear Lie algebra, $\mathfrak{gl}_{n}(\mathbb{C})$, whenever $n\geq d$. Due to its importance, Schur–Weyl duality has been generalized to a wide variety of settings. Relevant to this paper, these generalizations include when $\mathfrak{gl}_{n}(\mathbb{C})$ is replaced with affine and toroidal Lie algebras, or with the type $Q$ Lie superalgebra Flicker [2021]; Chari et al. [2024]; Sergeev [1984]. The symmetric group is then replaced with the affine symmetric group or Sergeev superalgebra, respectively. The purpose of the present paper is to provide a uniform proof that includes all of these, along with many additional cases of interest.
@@ 1.2. Main Result
For $d\geq 1$, let $S_{d}\wr A$ denote the wreath product superalgebra of $\mathds{k}S_{d}$ and $A$. Wreath product superalgebras naturally generalize many of the nice features of the symmetric groups and have long been of interest. On the Lie algebra side, for $n\geq 1$, let $\mathfrak{gl}_{n}(A)$ denote the Lie superalgebra over $\mathds{k}$ consisting of $n\times n$-matrices with entries from $A$ and Lie bracket given by the supercommutator. This also includes many examples of interest; see Section 2.3. We refer the reader to Sections 2.2 and 2.5 for details on the wreath product and Lie superalgebras.
@@
In Definition 2.4.1 we define a full subcategory, $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$, of $U(\mathfrak{gl}_{n}(A))$-supermodules consisting of supermodules which are polynomial representations of degree $d$ when restricted to the subalgebra $\mathfrak{gl}_{n}(\mathds{k})\subseteq\mathfrak{gl}_{n}(A)$ and also satisfy a certain condition that examples show is necessary. The image of $F^{A}_{n,d}$ is easily seen to land in $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$. Our main theorem is the following result.
@@
When $A=\mathds{k}[t_{1}^{\pm 1},\dotsc,t_{r}^{\pm 1}]$ or $A$ is the Clifford superalgebra on one generator, then we essentially recover the aforementioned generalizations of classical Schur–Weyl duality. Our result shows that we can further generalize to an arbitrary associative $\mathds{k}$-superalgebra. We also note that Example 3.8.1 shows that the bound $n\geq d+1$ is sharp. Thus, in some sense, this is the most general result along these lines that one could hope for.
Our general approach is similar to that of Flicker [2021]; Chari et al. [2024]: we prove the equivalence by bootstrapping from the classical Schur–Weyl equivalence. In comparing our arguments to theirs we find an instance of George Pólya’s Inventor’s Paradox: it turns out that our more general situation makes for easier proofs. As far as we are aware, this bootstrapping technique goes back to Chari–Pressley Chari and Pressley [1996].
@@ 1.3. Future Directions
There is a Schur–Weyl equivalence between the Lie superalgebra $\mathfrak{gl}_{m|n}(\mathbb{C})$ and the symmetric group Berele and Regev [1987]; Sergeev [1984]. We fully expect the methods used here should generalize to this case and yield an equivalence of categories between certain $\mathfrak{gl}_{m|n}(A)$-supermodules and $S_{d}\wr A$-supermodules. See Flicker [2023] where something like this is done for $A=\mathds{k}[t^{\pm 1}]$.
Classical Schur–Weyl duality holds for any infinite field, and one could hope to prove a positive characteristic analogue of the main theorem. Properly formulating a definition of the category $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$ will likely require working with an appropriate integral form for $U(\mathfrak{gl}_{n}(A))$. The existence of integral forms for the universal enveloping superalgebras of affine and map Lie (super)algebras suggests this should be possible Bagci and Chamberlin [2014]; Chamberlin [2013]; Garland [1978]. The representation theory of quantized and/or affinized wreath product superalgebras and their associated quantized Schur algebras is likewise interesting. See Lai et al. [2024]; Rosso and Savage [2020]; Savage [2020]; Wan and Wang [2008] and the references therein. It would be interesting to obtain a quantized and/or affinized version of the main theorem. See Chari and Pressley [1996] for the quantized affine case.
@@ 2.1. Notation and Conventions
$I(n,d)=\left\{\mathbf{i}=(i_{1},\dotsc,i_{d})\mid i_{1},\dotsc,i_{d}\in[1,n]\right\}$
denote the set of $d$-tuples with entries from $[1,n]$. Let $\tilde{I}(n,d)\subseteq I(n,d)$ be the subset of tuples that have pairwise distinct entries. Let
@@ 2.2. Lie Superalgebras
The set of matrices $\{E_{r,s}^{b}\mid r,s\in[1,n]\text{ and }b\in\mathcal{B}\}$ form a homogeneous $\mathds{k}$-basis for $\mathfrak{gl}_{n}(A)$. We view $\mathfrak{gl}_{n}(A)$ as a Lie superalgebra over $\mathds{k}$ via the supercommutator:
$E_{r_{1},s_{1}}^{a_{1}},E_{r_{2},s_{2}}^{a_{2}}]=\delta_{s_{1},r_{2}}E_{r_{1},s_{2}}^{a_{1}a_{2}}-(-1)^{\overline{a_{1}}\cdot\overline{a_{2}}}\delta_{s_{2},r_{1}}E_{r_{2},s_{1}}^{a_{2}a_{1}},$ (2.2.1)
for all $r_{1},s_{1},r_{2},s_{2}\in[1,n]$ and homogeneous $a_{1},a_{2}\in A$. As we did here, we often only give definitions and formulas on homogeneous elements. The general case is obtained by extending linearly.
@@
For short, write $h_{i}=E^{1}_{i,i}$ for $i=1,\dotsc,n$, and let $\mathfrak{h}$ denote the Cartan subalgebra of $\mathfrak{gl}_{n}(\mathds{k})$ with basis $h_{1},\dotsc,h_{n}$. Let $\varepsilon_{i}\in\mathfrak{h}^{*}=\operatorname{Hom}_{\mathds{k}}(\mathfrak{h},\mathds{k})$ be the linear functional determined by $\varepsilon_{i}(h_{j})=\delta_{i,j}$. Let $\Lambda(n)=\oplus_{i=i}^{n}\mathbb{Z}_{\geq 0}\varepsilon_{i}$. For $d\geq 0$, let
[…]
For short, given $\mathbf{i}=(i_{1},\dotsc,i_{d})\in I(n,d)$, write $\boldsymbol{\varepsilon}_{\mathbf{i}}=\sum_{k=1}^{d}\varepsilon_{i_{k}}\in\Lambda(n,d)$.
@@ 2.3. Examples
The above setup encompasses a number of cases of interest. We give a few examples.
If $A=\mathds{k}$, then $\mathfrak{gl}_{n}(\mathds{k})$ is the general linear Lie algebra. If $A=\operatorname{Cliff}:=\mathds{k}[c]/(c^{2}-1)$ is the Clifford superalgebra on one odd generator, then $\mathfrak{gl}_{n}(A)\cong\mathfrak{q}_{n}(\mathds{k})$ is the type Q Lie superalgebra. If $\Lambda=\mathds{k}[x]/(x^{2})$ is the Grassmann algebra on one odd generator, then the Lie superalgebra $\mathfrak{sl}_{2}(\Lambda)$ and its representations appear in the context of annular Khovanov homology Grigsby et al. [2018].
If $A=\mathds{k}[t_{1},\dotsc,t_{k}]$ or $A=\mathds{k}[t_{1}^{\pm 1},\dotsc,t_{k}^{\pm 1}]$, then $\mathfrak{gl}_{n}(A)$ is the general linear (multi)current Lie algebra or (multi)loop Lie algebra, respectively. The representations of (multi)loop Lie algebras are closely related to the level zero representations of the general linear affine and toroidal Lie algebras. Current algebras and their representations are central to the study of Yangians and, more recently, have appeared in the trace decategorifications of categorified quantum groups Molev [2007]; Beliakova et al. [2017]; Shan et al. [2017]. Generalizing the loop and current Lie algebras, if $A$ is a finitely-generated commutative $\mathds{k}$-algebra, then $\mathfrak{gl}_{n}(A)$ is a so-called map Lie algebra and its representations are an area of active research. A noncommutative example of this kind is when $A=\mathds{k}_{q}$, a quantum torus. In this case, $\mathfrak{sl}_{n}(\mathds{k}_{q})$ appears as the core of the elliptic quasi-simple Lie algebras in type A Berman et al. [1996].
Given a superalgebra $A$ and a superalgebra involution $\iota:A\to A$, one can define an associated superalgebra $\operatorname{Cliff}(A)$, which equals $A\otimes\operatorname{Cliff}$ as a superspace and has product given by $(a_{1}\otimes c^{k_{1}})(a_{2}\otimes c^{k_{2}})=(-1)^{\overline{a_{2}}\cdot\overline{k_{1}}}a_{1}\iota^{k_{1}}(a_{2})\otimes c^{k_{1}+k_{2}}$ for homogeneous $a_{1},a_{2}\in A$ and $k_{1},k_{2}\in\{0,1\}$. Then one could define $\mathfrak{q}_{n}(A)$ as $\mathfrak{gl}_{n}\left(A\otimes\operatorname{Cliff}\right)$. For example, taking $A=\mathds{k}$ and $\iota$ to be the identity map recovers the case $\mathfrak{q}_{n}(\mathds{k})\cong\mathfrak{gl}_{n}(\operatorname{Cliff})$ mentioned previously. Taking $A=\mathds{k}[t]$ or $A=\mathds{k}[t^{\pm 1}]$ and $\iota:A\to A$ determined by $t\mapsto-t$, or $A=\mathds{k}[t^{\pm 1}]$ and $\iota:A\to A$ determined by $t\mapsto t^{-1}$ recovers the twisted Lie superalgebras of type Q studied by Nazarov Nazarov [1999] and Chen–Guay Chen and Guay [2012]. The Lie superalgebra $\mathfrak{gl}_{n}(A)$ with $A=\mathds{k}\langle u^{\pm 1},c\rangle/(uu^{-1}=u^{-1}u=1,c^{2}=u,cu=uc)$ with $u$ even and $c$ odd is closely related to the loop algebra for the twisted affine type Q Lie superalgebra $\mathfrak{q}(n)^{(2)}$ studied in, for example, Hoyt and Serganova [2007]; Gorelik and Serganova [2008].
@@ 2.4. Representations
A $\mathfrak{gl}_{n}(A)$-supermodule is a $\mathds{k}$-superspace $M=M_{\bar{0}}\oplus M_{\bar{1}}$ which is a $U(\mathfrak{gl}_{n}(A))$-module where the action respects the parity in that $x.m\in M_{\overline{x}+\overline{m}}$ for all homogeneous $x\in U(\mathfrak{gl}_{n}(A))$ and $m\in M$. There is a natural $\mathbb{Z}_{2}$-grading on $\operatorname{Hom}_{\mathds{k}}(M,N)$ where $f:M\to N$ is declared to have parity $r\in\mathbb{Z}_{2}$ if $f(M_{s})\subseteq N_{s+r}$ for all $s\in\mathbb{Z}_{2}$. A homogeneous $\mathfrak{gl}_{n}(A)$-supermodule homomorphism is a homogeneous map $f:M\to N$ which satisfies $f(x.m)=(-1)^{\overline{f}\cdot\overline{x}}x.f(m)$ for all homogeneous $x\in U(\mathfrak{gl}_{n}(A))$ and $m\in M$. A general $\mathfrak{gl}_{n}(A)$-supermodule homomorphism is then a $\mathds{k}$-linear map $f:M\to N$ that can be written $f=f_{\bar{0}}+f_{\bar{1}}$ where $f_{r}$ is a homogeneous $\mathfrak{gl}_{n}(A)$-supermodule homomorphism of parity $r\in\mathbb{Z}_{2}$. In particular, note that we do not assume supermodule homomorphisms preserve parity.
@@
for all $i=1,\dotsc,n$. We write $M_{\lambda}$ for the subspace of all weight vectors of weight $\lambda$, and call $M$ a weight module if $M=\oplus_{\lambda\in\mathfrak{h}^{*}}M_{\lambda}$. The supercommutator formula (2.2.1) implies that
@@
Set $V_{n}=A^{\oplus n}$ to be column vectors of height $n$ with entries from $A$. This is naturally a $\mathfrak{gl}_{n}(A)$-supermodule with $\mathbb{Z}_{2}$-grading given by $V_{n,r}=A_{r}^{\oplus n}$ for $r\in\mathbb{Z}_{2}$, and with action given by matrix multiplication. Since $U(\mathfrak{gl}_{n}(A))$ is a Hopf superalgebra there is a natural action on $V_{n}^{\otimes d}$ for all $d\geq 1$. We follow the convention that $V_{n}^{\otimes 0}=\mathds{k}$ is the trivial supermodule.
For $i\in[1,n]$ and homogeneous $a\in A$, write $v_{i}^{a}$ for a column vector with $a$ in the $i$th row, and zeros elsewhere. The parity of $v_{i}^{a}$ is $\overline{a}$. Using this notation,
@@
is a homogeneous $\mathds{k}$-basis for $V_{n}^{\otimes d}$. From this it is easy to see that $V_{n}^{\otimes d}$ is a weight module, $v_{i_{1}}^{b_{1}}\otimes\dotsb\otimes v_{i_{d}}^{b_{d}}\in\left(V_{n}^{\otimes d}\right)_{\boldsymbol{\varepsilon}_{\mathbf{i}}}$, and that the weights of $V_{n}^{\otimes d}$ lie in $\Lambda(n,d)$.
Write $E_{n}=\mathds{k}^{\oplus n}$ for the natural representation of $\mathfrak{gl}_{n}(\mathds{k})$; that is, $E_{n}$ consists of column vectors of height $n$ with entries from $\mathds{k}$ where $\mathfrak{gl}_{n}(\mathds{k})$ acts by matrix multiplication. Recall that, for $d\geq 0$, the irreducible polynomial representations of degree $d$ of $\mathfrak{gl}_{n}(\mathds{k})$ are, by definition, the irreducible representations which appear as a composition factor of $E_{n}^{\otimes d}$.
For the purposes of this paper, a polynomial representation of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ will mean a $U(\mathfrak{gl}_{n}(\mathds{k}))$-supermodule which is isomorphic to a (possibly infinite) direct sum of irreducible polynomial representations of degree $d$. In particular, the category of polynomial representations of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$ is a semisimple category. If $M$ is a polynomial representation of degree $d$ for $U(\mathfrak{gl}_{n}(\mathds{k}))$, then $M$ is a weight module with
$M=\bigoplus_{\lambda\in\Lambda(n,d)}M_{\lambda}.$
@@
Remark 2.4.2.
Readers familiar with classical Schur–Weyl duality will not be surprised by the first condition in Definition 2.4.1. The second condition already appears in the study of polynomial representations of $\mathfrak{q}_{n}(\mathds{k})=\mathfrak{gl}_{n}(\operatorname{Cliff})$ [Grantcharov et al., 2010, Definition 1.5]. Furthermore, if $N$ is a $S_{d}\wr A$-supermodule, then the $U(\mathfrak{gl}_{n}(A))$-supermodule $V_{n}^{\otimes d}\otimes_{S_{d}\wr A}N$ satisfies both conditions. That is, the image of the functor $F_{n,d}^{A}$ lies in the category $\mathcal{P}_{d}(\mathfrak{gl}_{n}(A))$.
The reader may wonder if the second condition is either implied by the first condition or is otherwise an artifact of our argument. The next example shows that the second condition does not follow from being a polynomial representation of $U(\mathfrak{gl}_{n}(\mathds{k}))$; Example 3.8.1 gives an example of where the supermodule satisfies the first condition but not the second, and the supermodule also fails to be in the image of the functor $F_{n,d}^{A}$. From this point of view, our main result demonstrates that Definition 2.4.1 is an intrinsic characterization of the $\mathfrak{gl}_{n}(A)$-supermodules which are in the image of the functor $F_{n,d}^{A}$.